# References

### v1

@@ Acknowledgements
This work was done while visiting the Academia Sinica, Taipei for the Pacific-Rim Graph Theory Group Workshop in 2024 organized by Bruce Reed. Tony Huynh thanks Stefan Weltge from whom he learned about the tea sharing problem at the 26th Combinatorial Optimization Workshop in Aussois, France in 2024. We are also grateful to Nina Gantert and Timo Vilkas for providing valuable references and background information.
J. Pascal Gollin was supported by the Insitute for Basic Science (IBS-R029-Y3). Kevin Hendrey and Sang-il Oum were supported by the Insitute for Basic Science (IBS-R029-C1). Hao Huang was supported in part by a start-up grant at NUS and an MOE Academic Research Fund (AcRF) Tier 1 grant of Singapore. Tony Huynh was partially supported by Progetti Grandi di Ateneo Programme, Grant RG123188B3F7414A (ASGARD). Bojan Mohar was supported through the ERC Synergy grant KARST (European Union, ERC, KARST, project number 101071836). Wei-Hsuan Yu was supported by MOST of Taiwan under Grant109-2628-M-008-002-MY4. Xuding Zhu was supported by National Science Foundation of China with grant numbers NSFC 12371359 and U20A2068.
[…]
Appendix A Proposition 2.1 with general sharing moves
We present a direct proof of Corollary 1.2 (without needing to take a limit). This proof is adapted from our proof of Proposition 2.1.
An integer interval is a set of the form ${[a,b]}$, for some integers $a$ and $b$. Let $G$ be a graph and let ${r\in V(G)}$. Note that if $T$ is a finite nonempty subset of ${V(G)}$ which induces a connected subgraph, then the set ${\{d_{G}(r,v)\mid v\in T\}}$ is an integer interval. We start our first proof by an inequality on integer intervals.
@@
Since ${\frac{1+x}{x}}$ is a decreasing function for ${x>0}$, we may assume that ${x_{i}=x_{1}+i-1}$ by increasing values of $x_{i}$ whenever possible, as long as ${\{x_{1},x_{2},\ldots,x_{s},y_{1},y_{2},\ldots,y_{t}\}}$ is an integer interval. Therefore,
@@
Since ${\frac{y}{1+y}}$ is an increasing function for ${y>0}$, we may assume that ${y_{i}=x_{1}-i}$ by decreasing values of $y_{i}$ whenever possible, as long as ${\{x_{1},x_{2},\ldots,x_{s},y_{1},y_{2},\ldots,y_{t}\}}$ is an integer interval. Then
@@
We are now ready to prove the following strengthened version of Proposition 2.1.
@@
where the last inequality follows from Lemma A.1. Thus, $w$ is also $r$-feasible. ∎
Appendix B Lemma 3.4 with general sharing moves
We now present a stronger variant of Lemma 3.4 with general sharing moves. This will allow us to prove Proposition 3.5 without needing to take a limit. Throughout this appendix, $G$ is a fixed finite connected graph with $n$ vertices and $v$ is a fixed vertex of $G$. We use the same notation as in Section 3.
@@
We are now ready to prove the following strengthened version of Lemma 3.4.
@@
Let ${\sigma=(v_{1},\ldots,v_{n-1})}$ be an enumeration of $V(G)\smallsetminus\{v\}$ such that ${w^{\prime}(v_{i})\leq w^{\prime}(v_{i+1})}$ for all ${i\in[n-2]}$. By Lemma 3.3, ${\phi_{w^{\prime}}(w^{\prime}(v))=\phi_{\ell,w^{\prime}}(\sigma,w^{\prime}(v))}$, where
@@
It suffices to show that ${f_{(v_{s},v_{s+1},\ldots,v_{s+t-1});w^{\prime}}(a)\leq f_{(v_{s},v_{s+1},\ldots,v_{s+t-1});w}(a)}$. For each $j\in[t]$, let $d_{j}\coloneqq d_{G}(v,v_{s+j-1})$ and $w_{j}\coloneqq w(v_{s+j-1})$. Then, by Lemma B.1, we have
@@
So the only case left to consider is when ${v\in T}$. Since ${w^{\prime}(x)=\frac{w(T)}{{\left\lvert{T}\right\rvert}}}$ for every vertex $x\in T$, we may assume that ${T=\{v,v_{s},v_{s+1},\ldots,v_{s+t-2}\}}$ for some ${s\in[n-t+1]}$ with $t={\left\lvert{T}\right\rvert}$. We may also assume that ${w(v_{s+j-1})\leq w(v_{s+j})}$ for all $j\in[t-2]$. By Lemma 3.3,
@@
Since $\phi_{s+t-1,w}$ is strictly increasing, we may assume that $w(v)<w^{\prime}(v)$, because otherwise $\phi_{w^{\prime}}(w^{\prime}(v))=\phi_{s+t-1,w}(\sigma,w^{\prime}(v))\leq\phi_{s+t-1,w}(\sigma,w(v))\leq\phi_{w}(w(v))$. For each $j\in[t-1]$, let $d_{j}\coloneqq d_{G}(v,v_{s+j-1})$ and $w_{j}\coloneqq w(v_{s+j-1})$. Again, by Lemma B.1, we have
@@
This implies that $f_{(v_{s},v_{s+1},\ldots,v_{s+t-2});w}(w(v))\geq w^{\prime}(v)=f_{(v_{s},v_{s+1},\ldots,v_{s+t-2});w^{\prime}}(w^{\prime}(v))$ and therefore $\phi_{w}(w(v))\geq\phi_{s,w}(\sigma,w(v))\geq\phi_{s,w^{\prime}}(\sigma,w^{\prime}(v))=\phi_{w^{\prime}}(w^{\prime}(v))$, where the last equality follows from Lemma 3.3. ∎
Combining Lemma B.2 with Lemma 3.1Lemma 3.2, and Lemma 3.3 gives the desired proof of Proposition 3.5.

### v2

# Sharing tea on a graph

@@ Acknowledgements
This work was done while visiting the Academia Sinica, Taipei for the Pacific Rim Graph Theory Group Workshop in 2024 organized by Bruce Reed. Tony Huynh thanks Stefan Weltge from whom he learned about the tea sharing problem at the 26th Combinatorial Optimization Workshop in Aussois, France in 2024. We are also grateful to Nina Gantert and Timo Vilkas for providing valuable references and background information.
J. Pascal Gollin was supported by the Institute for Basic Science (IBS-R029-Y3). Kevin Hendrey, Tony Huynh, and Sang-il Oum were supported by the Institute for Basic Science (IBS-R029-C1). Hao Huang was supported in part by a start-up grant at NUS and an MOE Academic Research Fund (AcRF) Tier 1 grant of Singapore. Bojan Mohar was supported through the ERC Synergy grant KARST (European Union, ERC, KARST, project number 101071836). Wei-Hsuan Yu was supported by MOST of Taiwan under Grant109-2628-M-008-002-MY4. Xuding Zhu was supported by National Science Foundation of China with grant numbers NSFC 12371359 and U20A2068.
Appendix A Proposition˜2.1 with general sharing moves
We now present a direct proof of Corollary˜1.2. This proof does not use Theorem˜1.1, nor does it require taking a limit (as in the proof of Corollary˜1.2 given in Section˜1). This proof is adapted from our proof of Proposition˜2.1.
An integer interval is a set of the form [a,b]{[a,b]}, for some integers $a$ and $b$. Let $G$ be a graph and let $r\in V(G)$}. Note that if $T$ is a finite nonempty subset of V(G){V(G)} which induces a connected subgraph, then the set $\{d_{G}(r,v)\mid v\in T\}$} is an integer interval. We start our first proof by an inequality on integer intervals.
@@
Since $\frac{1+x}{x}$} is a decreasing function for x>0{x>0}, we may assume that $x_{i}=x_{1}+i-1$} by increasing values of $x_{i}$ whenever possible, as long as $\{x_{1},x_{2},\ldots,x_{s},y_{1},y_{2},\ldots,y_{t}\}$} is an integer interval. Therefore,
@@
Since $\frac{y}{1+y}$} is an increasing function for y>0{y>0}, we may assume that $y_{i}=x_{1}-i$} by decreasing values of $y_{i}$ whenever possible, as long as $\{x_{1},x_{2},\ldots,x_{s},y_{1},y_{2},\ldots,y_{t}\}$} is an integer interval. Then
@@
We are now ready to prove the following strengthened version of Proposition˜2.1.
@@
where the last inequality follows from Lemma˜A.1. Thus, $w$ is also $r$-feasible. ∎
Appendix B Lemma˜3.4 with general sharing moves
We now present a stronger variant of Lemma˜3.4 with general sharing moves. This will allow us to prove Proposition˜3.5 without needing to take a limit. Throughout this appendix, $G$ is a fixed finite connected graph with $n$ vertices and $v$ is a fixed vertex of $G$. We use the same notation as in Section˜3.
@@
We are now ready to prove the following strengthened version of Lemma˜3.4.
@@
Let $\sigma=(v_{1},\ldots,v_{n-1})$} be an enumeration of $V(G)\smallsetminus\{v\}$ such that $w^{\prime}(v_{i})\leq w^{\prime}(v_{i+1})$} for all $i\in[n-2]$}. By Lemma˜3.3, $\phi_{w^{\prime}}(w^{\prime}(v))=\phi_{\ell,w^{\prime}}(\sigma,w^{\prime}(v))$}, where
@@
It suffices to show that $f_{(v_{s},v_{s+1},\ldots,v_{s+t-1});w^{\prime}}(a)\leq f_{(v_{s},v_{s+1},\ldots,v_{s+t-1});w}(a)$}. For each $j\in[t]$, let $d_{j}\coloneqq d_{G}(v,v_{s+j-1})$ and $w_{j}\coloneqq w(v_{s+j-1})$. Then, by Lemma˜B.1, we have
@@
So the only case left to consider is when $v\in T$}. Since $w^{\prime}(x)=\frac{w(T)}{{\left\lvert{T}\right\rvert}}$} for every vertex $x\in T$, we may assume that $T=\{v,v_{s},v_{s+1},\ldots,v_{s+t-2}\}$} for some $s\in[n-t+1]$} with $t={\left\lvert{T}\right\rvert}$. We may also assume that $w(v_{s+j-1})\leq w(v_{s+j})$} for all $j\in[t-2]$. By Lemma˜3.3,
@@
Since $\phi_{s+t-1,w}$ is strictly increasing, we may assume that $w(v)<w^{\prime}(v)$, because otherwise $\phi_{w^{\prime}}(w^{\prime}(v))=\phi_{s+t-1,w}(\sigma,w^{\prime}(v))\leq\phi_{s+t-1,w}(\sigma,w(v))\leq\phi_{w}(w(v))$. For each $j\in[t-1]$, let $d_{j}\coloneqq d_{G}(v,v_{s+j-1})$ and $w_{j}\coloneqq w(v_{s+j-1})$. Again, by Lemma˜B.1, we have
@@
This implies that $f_{(v_{s},v_{s+1},\ldots,v_{s+t-2});w}(w(v))\geq w^{\prime}(v)=f_{(v_{s},v_{s+1},\ldots,v_{s+t-2});w^{\prime}}(w^{\prime}(v))$ and therefore $\phi_{w}(w(v))\geq\phi_{s,w}(\sigma,w(v))\geq\phi_{s,w^{\prime}}(\sigma,w^{\prime}(v))=\phi_{w^{\prime}}(w^{\prime}(v))$, where the last equality follows from Lemma˜3.3. ∎
Combining Lemma˜B.2 with Lemma˜3.1, Lemma˜3.2, and Lemma˜3.3 gives the desired proof of Proposition˜3.5.